Three audience members are chosen at random to participate in a magic show. If there are 75 people in attendance, What is the probability you and your two friends are chosen?


1
Expert's answer
2021-01-07T14:07:18-0500

"P=\frac{1}{75}*\frac{1}{74}*\frac{1}{73}*P_3^3=\frac{1}{75}*\frac{1}{74}*\frac{1}{73}*6=\\=\frac{1}{67525}=8.89*10^{-5}" .